# ArmGS: Composite Gaussian Appearance Refinement for Modeling Dynamic Urban Environments

### v1

@@ 4.1 Datasets and Evaluation Protocol
We compare our ArmGS approach with several state-of-the-art methods which can be categorized into two types: (1) 3DGS-based methods, including StreetGaussians [28], S3Gaussian [11], 3DGS [12] and HUGS [34]; (2) NeRF-based methods, including EmerNeRF [29], ML-NSG [7], SUDS [24], NSG [19], StreetSurf [9] and MARS [27]. Please refer to the supplementary material for the implementation details of our approach.
4.2 Evaluation on Waymo
@@
We present the quantitative results on Waymo in Tab. 1. We can see that our approach achieves significantly better performance compared with the state-of-the-arts in both the novel view synthesis and the image reconstruction settings. Specifically, for image reconstruction, our approach achieves the best PSNR of 38.1 dB, SSIM of 0.957 and LPIPS of 0.064, outperforming the state-of-the-art competitors, such as StreetGaussians and EmerNeRF. For novel view synthesis, our approach still achieves the best results compared with the state-of-the-art methods. For instance, our approach achieves PSNR of 35.7 dB, which significantly surpass the second-best StreetGaussians by 1.1 dB. Furthermore, in Fig. 3, we present some qualitative results of scene modeling under foggy, sunny, cloudy and rainy conditions on Waymo. It can be seen that our approach renders more fine-grained details of the background scenes and dynamic actors. For example, as shown in the first row, most methods fail to render the appearance change of traffic lights, while our approach render this fine-grained variation; and as shown in the second row, our approach is able to model appearance of the distant vehicles, while other methods render results with artifacts. Moreover, our experimental results show that the rendering speeds of the compared NeRF-based methods, e.g., EmerNeRF, are usually less than 1 FPS; in comparison, the 3DGS-based methods (including our approach) are able to achieve real-time rendering ($>$30FPS).
4.3 Evaluation on KITTI
@@
We report the quantitative and qualitative results on the KITTI dataset in Tab. 1 and Fig. 4, respectively. Overall, our approach achieves significantly better performance compared with the state-of-the-art methods. Specifically, from Tab. 1, we can see that our approach achieves the best PSNR of 30.3 dB, SSIM of 0.931 and LPIPS of 0.036, significantly outperforming the state-of-the-art methods. Furthermore, from Fig. 4, we can see that compared with StreetGaussians, our approach is able to render better image quality with more fine-grained details, such as vehicles highlighted in the first row and the lane line highlighted in the second row.
4.4 Evaluation on NOTR
To further probe the upper bound reconstruction capability of our approach, we follow [29] to conduct the scene reconstruction experiment on the NOTR dataset [29]. We report the quantitative and qualitative results in Tab. 2 and Fig. 5, respectively. As shown in Tab. 2, compared with the state-of-the-art StreetSurf [9], EmerNeRF [29] and StreetGaussians [28], our approach achieves better performance on both the static-32 and dynamic-32 splits. From Fig. 5, we can see that our approach is able to model fine-grained details of driving scenes across cameras; in comparison, StreetGaussians renders driving scenes with more artifacts, such as blurred vehicles. These results indicate that our approach is able to faithfully model complex driving scenes and render high-quality results.
@@ 4.4 Evaluation on NOTR
Table 2: Results on the NOTR dataset. PSNRs are reported.
@@
Table 3: Results on the VKITTI2 dataset.
4.5 Evaluation on VKITTI2
Our approach is not limited to modeling real-world driving scenes, but can also be used to model synthetic driving scenes. To show the performance of our approach for synthetic driving scenes modeling, we follow [34] conducting experiments on the VKITTI2 dataset [1]. We report the results in Tab. 3. Although synthetic scenes usually have fewer appearance changes of scenes and objects across frames, our approach still achieves better results compared with the state-of-the-art HUGS, demonstrating the superiority of our approach. These results indicate that our approach is also applicable to synthetic driving scene modeling.
4.6 Ablation Study
Dynamic Actor Modeling Effectiveness Analysis.
To evaluate the effectiveness of our dynamic actor level refinement, we follow [28] projecting the 3D bounding box onto the 2D image plane to obtain the mask regions of moving objects and compute PSNR within the mask regions. This mitigates the effect of background on dynamic actors when evaluating dynamic actors. As shown in Tab. 4, our approach achieves better results for the rendered dynamic actor regions compared with StreetGaussians and EmerNeRF. When using class-wise actor encoding, our approach achieves similar performance, but when dynamic actor level refinement is not used, the PSNRs of our approach are reduced by 0.4 dB and 0.2 dB in the reconstruction and novel view synthesis setting, respectively. When replacing our dynamic actor level refinement design with the 4DGS HexPlane [26], the results are still worse than our design.
@@ 4.6 Ablation Study
Ours w/ 4DGS HexPlane [26] 30.6 35.7
EmerNeRF [29] 21.7 26.8
StreetGaussians [28] 30.2 34.8
Table 4: Effectiveness analysis of dynamic actor modeling on the Waymo dataset. PSNRs of the moving object regions are reported.
@@
Ours w/ HUGS appearance [34] 34.8 36.5
Table 5: Effectiveness analysis of local and global appearance refinement on the Waymo Dataset. Results are in terms of PSNR.
Local and Global Appearance Refinement Effectiveness Analysis.
In Tab. 5, we evaluate the effectiveness of the local composite Gaussian level refinement and the global image level refinement. It can be seen that our approach with both local and global appearance refinement performs the best, while without using local level refinement or global level refinement, the performance of our approach degrades. Besides, using HUGS appearance modeling in lieu of our design results in worse performance. These results verify the effectiveness of our design.
5 Conclusion
This work presents a composite Gaussian appearance refinement approach (ArmGS) to urban environments modeling for autonomous driving simulation. The key idea is to learn a set of transformation parameters to refine 3D Gaussians from multi-level granularities, ranging from local Gaussian level to global image level and dynamic actor level. Our thorough experiments on four autonomous driving datasets demonstrate that our approach is capable of rendering more fine-grained details of driving scenes across frames and camera viewpoints, achieving superior performance compared with the state-of-the-art methods.
Supplementary Material
In this supplementary material, we present implementation details of our approach in Sec. 6, show more experimental analysis and discussion in Sec. 7, and discuss the limitation and future work of our approach in Sec. 8.
6 Implementation Details
We implement our approach with Python and PyTorch. We initialize driving scene Gaussians with LiDAR point clouds and SfM points extracted by COLMAP [22], and train our model with 30000 iterations using the Adam [13] optimizer. We set the initial position learning rate to $1.6e^{-4}$ and decay to $1.6e^{-6}$, while the learning rates for rotation, scaling, opacity and spherical harmonics are set to $1e^{-3}$, $5e^{-3}$, $5e^{-2}$ and $2.5e^{-3}$, respectively. We construct the local affine transformation learner with three linear layers, the global affine transformation learner with four linear layers and the class encoder with two linear layers, where ReLU activations are used between layers. For the spatial-temporal encoder and the deformation head, we employ two linear layers with ReLU activations between layers. By default, we use $\mathcal{F}^{f}$ for dynamic actor position encoding. For training objective loss, we empirically set $\lambda_{1}{=}0.2$, $\lambda_{2}{=}0.01$, $\lambda_{3}{=}0.05$ and $\lambda_{4}{=}0.1$. Following [12], Gaussian splitting and merging for adaptive density control are performed every 100 iterations from 500 iterations and until 15000 iterations.
7 Further Analysis and Discussion
Novel View Synthesis with Lane Shift Evaluation.
We evaluate the performance of our approach for the extrapolated novel view synthesis by moving the ego-vehicle to left or right by 1, 2 and 3 meters. As shown in Tab. 6, our approach achieves lower FIDs than StreetGaussians under different lane shift distances. From Fig. 6, we can see that our approach produces fewer rendering artifacts for extrapolated novel views synthesis. These results indicate that our approach is capable of simulating extrapolated novel views with better quality.
Driving Scene Decomposition.
In Fig. 7, we show the results of our approach for driving scene decomposition of static background and dynamic actors. It can be seen that our approach can well separate static background and dynamic actors. This indicates that although background and dynamic actors are jointly optimized with the multi-level appearance modeling scheme in our approach, this scheme does not cause the background Gaussians and dynamic actor Gaussians to interfere with each other.
@@ 7 Further Analysis and Discussion
ArmGS (Ours) 73.1 104.5 128.1
StreetGaussians [28] 81.8 113.3 140.2
Table 6: Lane shift evaluation on the Waymo dataset.
Figure 6: Novel view synthesis with lane shift evaluation on the Waymo dataset.
Figure 7: Driving scene decomposition results of our approach.
Figure 8: Examples of driving scene editing with our rendering results. The first row shows rendered results by swapping the vehicles and the second row shows synthesizing a corner case of vehicles driving in the wrong lane.
Driving Scene Editing.
In Fig. 8, we show some examples of driving scene editing with our rendering results, including swapping positions of dynamic vehicles in a driving scene or synthesizing a corner case of vehicles driving in the wrong lane. From Fig. 8, we can see that these edited scenes are still photorealistic. These results show that our approach can potentially benefit other downstream autonomous driving simulation tasks, such as synthesizing challenging corner cases.
8 Limitation and Future Work
Despite the promising performance, our approach still has limitations. Specifically, our approach is limited to rendering images and lacks the ability to simulate LiDAR point clouds. How to simulate high-fidelity LiDAR point clouds with 3DGS for autonomous driving scene modeling remains an open question. Besides, incorporating our approach into an autonomous driving simulator for closed-loop evaluation is an interesting direction. This can help to test the safety and reliability of autonomous vehicles. Our future work aims to solve these problems and build a comprehensive autonomous driving simulator.

### v2

@@ IV-A Datasets and Evaluation Protocol
We compare our ArmGS approach with several state-of-the-art methods which can be categorized into two types: (1) 3DGS-based methods, including StreetGS [31], OmniRe [3], SplatFlow [26], S3Gaussian [13], 3DGS [14] and HUGS [37]; (2) NeRF-based methods, including EmerNeRF [32], ML-NSG [9], SUDS [27], NSG [21], StreetSurf [11] and MARS [30].
IV-B Implementation Details
We implement our approach with Python and PyTorch. We initialize driving scene Gaussians with LiDAR point clouds and SfM points extracted by COLMAP, and train our model with 30000 iterations using the Adam optimizer. We adopt the training and evaluation splits for each dataset as described in Sec. IV-A. We set the initial position learning rate to $1.6e^{-4}$ and decay it to $1.6e^{-6}$. The learning rates for rotation, scaling, opacity and spherical harmonics are set to $1e^{-3}$, $5e^{-3}$, $5e^{-2}$ and $2.5e^{-3}$, respectively. We construct the local affine transformation learner with three linear layers, the global affine transformation learner with four linear layers and the class encoder with two linear layers, where ReLU activations are used between layers. For the spatial-temporal encoder and the deformation head, we employ two linear layers with ReLU activations between layers. By default, we use $\mathcal{F}^{f}$ for dynamic actor position encoding. For training loss, we empirically set $\lambda_{1}{=}0.2$, $\lambda_{2}{=}0.01$, $\lambda_{3}{=}0.05$ and $\lambda_{4}{=}0.1$. Gaussian splitting and merging for adaptive density control are performed every 100 iterations from 500 iterations and until 15000 iterations.
Figure 4: Qualitative comparison with the state-of-the-art methods on KITTI. We highlight some fine-grained details, e.g., vehicles and lane lines.
IV-C Evaluation on Waymo
@@
We present the quantitative results on Waymo in Tab. I. We can see that our approach achieves significantly better performance compared with the state-of-the-arts in both the image reconstruction and the novel view synthesis settings. Specifically, for image reconstruction, our approach achieves the best PSNR of 38.1 dB, SSIM of 0.957 and LPIPS of 0.064, outperforming the state-of-the-art competitors, such as StreetGS, OmniRe and EmerNeRF. For novel view synthesis, our approach still achieves the best results compared with the state-of-the-art methods. For instance, our approach achieves PSNR of 35.7 dB, which significantly surpass the second-best StreetGS by 1.1 dB. Furthermore, in Fig. 3, we present some qualitative results of scene modeling under foggy, sunny, cloudy and rainy conditions on Waymo. It can be seen that our approach renders more fine-grained details of the background scenes and dynamic actors. For example, as shown in the first row, most methods fail to render the appearance change of traffic lights, while our approach render this fine-grained variation; and as shown in the second row, our approach is able to model appearance of the distant vehicles, while other methods render results with artifacts. Moreover, our experimental results show that the rendering speeds of the compared NeRF-based methods, e.g., EmerNeRF, are usually less than 1 FPS; in comparison, the 3DGS-based methods (including our approach) are able to achieve real-time rendering (>>30FPS).
IV-D Evaluation on KITTI
@@
We report the quantitative and qualitative results on the KITTI dataset in Tab. I and Fig. 4, respectively. Overall, our approach achieves significantly better performance compared with the state-of-the-art methods. Specifically, from Tab. I, we can see that our approach achieves the best PSNR of 30.3 dB, SSIM of 0.931 and LPIPS of 0.036, significantly outperforming the state-of-the-art methods. Furthermore, from Fig. 4, we can see that compared with StreetGS, our approach is able to render better image quality with more fine-grained details, such as vehicles highlighted in the first row and the lane line highlighted in the second row.
Figure 5: Qualitative comparison on NOTR. We highlight some fine-grained details, e.g., vehicles and pedestrians.
TABLE II: Results on NOTR. PSNRs are reported.
@@
IV-E Evaluation on NOTR
To further probe the upper bound reconstruction capability of our approach, we follow [32] to conduct the scene reconstruction experiment on the NOTR dataset [32]. We report the quantitative and qualitative results in Tab. II and Fig. 5, respectively. As shown in Tab. II, compared with the state-of-the-art StreetSurf [11], EmerNeRF [32] and StreetGS [31], our approach achieves better performance on both the static-32 and dynamic-32 splits. From Fig. 5, we can see that our approach is able to model fine-grained details of driving scenes across cameras; in comparison, StreetGS renders driving scenes with more artifacts, such as blurred vehicles. These results show that our approach can faithfully model complex driving scenes and render high-quality results.
Figure 6: Urban scene modeling under diverse conditions using our approach.
IV-F Evaluation on VKITTI2
Our approach is not limited to modeling real-world driving scenes, but can also be used to model synthetic driving scenes. To show the performance of our approach for synthetic driving scenes modeling, we follow [37] conducting experiments on the VKITTI2 dataset [1]. We report the results in Tab. III. Although synthetic scenes usually have fewer appearance changes of scenes and objects across frames, our approach still achieves better results compared with the state-of-the-art HUGS.
TABLE III: Results on VKITTI2.
@@ IV-F Evaluation on VKITTI2
IV-G Ablation Study
Dynamic Actor Modeling Effectiveness Analysis
To evaluate the effectiveness of our dynamic actor level refinement, we follow [31] projecting the 3D bounding box onto the 2D image plane to obtain the mask regions of moving objects and compute PSNR within the mask regions. This mitigates the effect of background on dynamic actors when evaluating dynamic actors. As shown in Tab. IV, our approach achieves better results for the rendered dynamic actor regions compared with StreetGS and EmerNeRF. When using class-wise actor encoding, our approach achieves similar performance. When the dynamic actor level refinement is removed, the performance of our approach degrades in both reconstruction and novel view synthesis settings. When replacing our design with the 4DGS HexPlane [29], the result becomes worse. We conjecture that, compared with our design, the 4DGS HexPlane requires structured factorization for spatial-temporal modeling which constrains its capability to capture variations of dynamic actors.
TABLE IV: Effectiveness analysis of dynamic actor modeling on Waymo. PSNRs of the moving object regions are reported.
@@ IV-G Ablation Study
Ours w/ 4DGS HexPlane 30.6 35.7
EmerNeRF 21.7 26.8
StreetGS 30.2 34.8
Local and Global Appearance Refinement Effectiveness Analysis
In Tab. V, we evaluate the effectiveness of the local composite Gaussian level refinement and the global image level refinement. It can be seen that our approach with both local and global appearance refinement performs the best, while without using local level refinement or global level refinement, the performance of our approach degrades. Besides, using HUGS appearance modeling in lieu of our design results in worse performance.
TABLE V: Effectiveness analysis of local and global appearance refinement on Waymo. Results are in terms of PSNR.
@@
Ours w/ HUGS Appearance 34.8 36.5
TABLE VI: Ablation analysis of more components.
@@
ArmGS 38.1
without LiDAR depth loss 37.9
without LiDAR initial points 37.5
without sky GS model and mask 37.6
without actor pose optimization 37.7
add camera pose optimization 38.1
Ablation Analysis of More Components
In Tab. VI, we report ablation analysis of more components. We can see that: (1) LiDAR depth loss has a minor impact on our approach; (2) Without LiDAR initial points, the performance of our approach only slightly decreases; (3) Sky Gaussians can slightly affect the model performance; (4) Actor pose optimization helps the model to achieve better performance; (5) Adding additional camera pose optimization to our approach does not bring performance improvement.
Urban Scene Modeling under Diverse Weather Conditions
Real-world urban scenes often involve diverse weather conditions. Although our approach is not specifically designed for urban scene modeling under extreme weather conditions, our approach still performs well for different urban scene conditions, e.g., foggy, sunny, cloudy, rainy and nighttime. Fig. 6 shows some results of our approach for urban scene modeling under diverse conditions.
V CONCLUSION
This work presents a composite Gaussian appearance refinement approach to urban environments modeling. The key idea is to learn a set of transformation parameters to refine 3D Gaussians from multi-level granularities, ranging from local Gaussian level to global image level and dynamic actor level. Our thorough experiments on four autonomous driving datasets demonstrate that our approach is capable of rendering more fine-grained details of driving scenes across frames and camera viewpoints, achieving superior performance compared with the state-of-the-art methods.